OpenStudy (anonymous):

How long will it take an investment to double in value if the interest rate is 8% compounded continuously?

OpenStudy (anonymous):

How long will it take an investment to double in value if the interest rate is 8% compounded continuously?

OpenStudy (kropot72):

\[A=Pe ^{rt}\] P = principal amount (initial investment) r = annual interest rate (as a decimal) t = number of years A = amount after time t \[\frac{A}{P}=e ^{rt}=e ^{0.08t}=2\] Now you need to solve for t.

OpenStudy (kropot72):

@kglaguardia Do you know how to solve for t?

OpenStudy (anonymous):

\[2=ln e ^{0.08t}\]...lets you bring down the exponent so it would be \[2=0.08t \ln e\]...bring over the 0.08 and lne is equal to 1

OpenStudy (kropot72):

\[e ^{0.08t}=2\] Taking natural logs of both sides gives: \[0.08t=\ln 2\] \[t=\frac{\ln 2}{0.08}\]

OpenStudy (anonymous):

right. forgot about the other side.